### 20 Mar 2026 v1

# Colourings of Cayley graphs of finite $3$-groups

Abstract.
Colouring problems arising from group-based constructions provide a natural link between combinatorics and algebra, particularly in the study of Cayley graphs and Latin squares. We introduce colouring bijections of finite groups, a class of permutations encoding proper vertex colourings of associated Cayley-type graphs, extending classical notions such as complete and strong complete mappings. We prove that every finite $3$-group without a cyclic maximal subgroup admits a colouring bijection. Consequently, for such groups $G$, the graph $\mathscr{G}_{3}(G)$ admits a proper colouring with $|G|$ colours. These results show that the existence of colouring bijections is governed by structural properties of $3$-groups, revealing a new connection between group theory and combinatorial colouring problems.
Key words and phrases:
Cayley graph, finite $3$-group, graph colouring, complete mapping, strong complete mapping
2020 Mathematics Subject Classification:
05C15, 20D15, 05C25
This work was supported by Grant WZ/WI-IIT/2/2025 from the Bialystok University of Technology and funded by the Ministry of Science and Higher Education of Poland through resources allocated for research.
Introduction and motivation
Colouring problems form a central theme in combinatorics, with numerous connections to graph theory, design theory, and algebraic structures. In particular, colourings arising from group-based constructions, such as Cayley graphs, provide a natural bridge between algebra and combinatorics.
In this paper, we study colouring bijections associated with finite $3$-groups and interpret them combinatorially via Cayley graphs. More precisely, we consider colourings of Cayley graphs induced by bijections on group elements and investigate how the algebraic structure of the group constrains the existence and properties of such colourings.
Recall that a bijection $\sigma\colon G\to G$ is called a complete mapping if the map $x\mapsto x\sigma(x)$ is also a bijection. Complete mappings arise naturally in the study of Latin squares and orthomorphisms, and their existence has been extensively investigated. The problem of determining which finite groups admit complete mappings culminated in the resolution of the Hall–Paige conjecture. The classical theorem of Hall and Paige [10], completed in later work by Wilcox [12], Evans [7, 9], and Bray, Cai, Cameron, Spiga, and Zhang [5], states that a finite group admits a complete mapping if and only if its Sylow $2$-subgroups are not cyclic of order greater than $1$.
The graphs considered in this paper arise as a higher-dimensional analogue of this construction. Following [3], for a finite group $G$ we define the Cayley graph
$\mathscr{G}_{3}(G)=\mathrm{Cay}(G^{3},\mathbf{S}_{3}),$
where
$\mathbf{S}_{3}=\{(g,e,e),(e,g,e),(e,e,g),(g,g,e),(e,g,g),(g,g,g)\mid g\in G\setminus\{e\}\}.$
The graphs $\mathscr{G}_{m}(G)$ were introduced and studied in [3], where it was shown that they determine the underlying group: for finite groups $G$ and $H$, the graphs $\mathscr{G}_{m}(G)$ and $\mathscr{G}_{m}(H)$ are isomorphic if and only if $G\cong H$. Each graph $\mathscr{G}_{3}(G)$ contains cliques of size $|G|$, and therefore
$\chi(\mathscr{G}_{3}(G))\geq|G|.$
To study the equality case, we introduce the following notion. A bijection $\sigma\colon G\to G$ is called a colouring bijection if the maps
$x\mapsto\sigma(x)x,\qquad x\mapsto x^{-1}\sigma(x),\qquad x\mapsto x^{-1}\sigma(x)x$
are bijections of $G$. The existence of such a bijection implies that $\mathscr{G}_{3}(G)$ admits a proper colouring with $|G|$ colours (see Proposition 1.1). Thus, colouring bijections may be viewed as a natural higher-dimensional analogue of complete mappings in the context of graph colourings arising from Latin squares. A group admitting a colouring bijection will be called colourable.
In the abelian case, colouring bijections coincide with strong complete mappings, that is, bijections $\sigma\colon G\to G$ such that the maps $x\mapsto x\sigma(x)$ and $x\mapsto x^{-1}\sigma(x)$ are bijections. In general, there is no canonical transformation converting strong complete mappings into colouring bijections, since the expressions $\sigma(x)x$, $x^{-1}\sigma(x)$, and $x^{-1}\sigma(x)x$ are not related by uniform compositions with inversion, translations, or conjugation. The existence of strong complete mappings for abelian $3$-groups was determined by Evans [8], and further constructions were obtained in [4].
For nonabelian $3$-groups, the existence problem for strong complete mappings was studied by Akhtar and Gagola [1]. Within this class, a major obstruction is provided by the family of groups of maximal class
$L_{r}=\langle a,b\mid a^{3^{r-1}}=b^{3}=1,\ bab^{-1}=a^{1+3^{r-2}}\rangle,\qquad r\geq 4.$
Akhtar and Gagola [1] proved that every noncyclic $3$-group not isomorphic to $L_{r}$ for $r\geq 4$ admits a strong complete mapping.
A similar phenomenon occurs for colouring bijections. We establish the following result.
Main result. Every noncyclic $3$-group not isomorphic to $L_{r}$ for $r\geq 4$ admits a colouring bijection.
Consequently,
$\chi(\mathscr{G}_{3}(G))=|G|$
for all such groups.
The proof combines structural properties of finite $3$-groups with a lifting approach. We first construct colouring bijections explicitly for nonabelian groups of order $27$ using computer-assisted search. These constructions serve as base cases for an inductive procedure extending colouring bijections along suitable normal subgroups.
A key step is to show that every noncyclic $3$-group contains a characteristic subgroup isomorphic to $C_{3}\times C_{3}$, unless it belongs to the exceptional family $L_{r}$. This reduction enables the lifting arguments to be applied in a uniform way.
The paper is organized as follows. Section 1 introduces notation and preliminary results on $3$-groups and colouring bijections. Section 2 briefly discusses colourability of $3$-groups of order at most $27$. Section 3 develops the lifting techniques and establishes the necessary extension results. Section 4 contains the proof of the main theorem. The appendix provides explicit constructions of colouring bijections for groups of order $27$ obtained by exhaustive computer search.
1. Preliminary results
We begin by proving the statement announced in the introduction, which motivates the study of colouring bijections.
To a map $\sigma\colon G\to G$ we associate three maps $\Delta^{(i)}_{\sigma}\colon G\to G$ defined by
$\Delta^{(1)}_{\sigma}(x)=\sigma(x)x,\qquad\Delta^{(2)}_{\sigma}(x)=x^{-1}\sigma(x),\qquad\Delta^{(3)}_{\sigma}(x)=x^{-1}\sigma(x)x.$
Thus, a bijection $\sigma$ is called a colouring bijection if all the maps $\Delta^{(i)}_{\sigma}$ are bijections.
Proposition 1.1.
Suppose that a group $G$ admits a colouring bijection $\sigma\colon G\to G$. Then the function $c\colon G^{3}\to G$ defined by
$c(x,y,z)=x^{-1}\sigma(y)z$
is an $|G|$-colouring of the graph $\mathscr{G}_{3}(G)$.
Proof.
Recall that $\mathscr{G}_{3}(G)=\operatorname{Cay}(G^{3},\mathbf{S}_{3})$, where
$\mathbf{S}_{3}=\{(g,e,e),(e,g,e),(e,e,g),(g,g,e),(e,g,g),(g,g,g)\mid g\in G\setminus\{e\}\}.$
Thus every neighbour of $(x,y,z)$ has one of the forms
(gx,y,z),(x,gy,z),(x,y,gz),(gx,gy,z),(x,gy,gz),(gx,gy,gz),(gx,y,z),\ (x,gy,z),\ (x,y,gz),\ (gx,gy,z),\ (x,gy,gz),\ (gx,gy,gz),
for some $g\neq e$.
We show that adjacent vertices receive distinct colours.
For the first three types of neighbours, we have
$c(gx,y,z)=x^{-1}g^{-1}\sigma(y)z,\qquad c(x,gy,z)=x^{-1}\sigma(gy)z,\qquad c(x,y,gz)=x^{-1}\sigma(y)gz,$
so none of these equals $c(x,y,z)=x^{-1}\sigma(y)z$, since $g\neq e$ and $\sigma$ is injective.
For the remaining three types, equality of colours is equivalent to
$c(gx,gy,z)=c(x,y,z)$ $\iff g^{-1}\sigma(gy)=\sigma(y)$
$\iff(gy)^{-1}\sigma(gy)=y^{-1}\sigma(y),$
$c(x,gy,gz)=c(x,y,z)$ $\iff\sigma(gy)g=\sigma(y)$
$\iff\sigma(gy)gy=\sigma(y)y,$
$c(gx,gy,gz)=c(x,y,z)$ $\iff g^{-1}\sigma(gy)g=\sigma(y)$
$\iff(gy)^{-1}\sigma(gy)gy=y^{-1}\sigma(y)y.$
These contradict injectivity of $\Delta_{\sigma}^{(2)}$, $\Delta_{\sigma}^{(1)}$, and $\Delta_{\sigma}^{(3)}$, respectively.
Hence adjacent vertices always receive distinct colours, so $c$ is an $|G|$-colouring of $\mathscr{G}_{3}(G)$. ∎
Since for a prime $p$ the map $x\mapsto x^{p}$ is a bijection in any group whose order is not divisible by $p$, we obtain the following corollary.
Corollary 1.2.
Every group whose order is divisible by neither $2$ nor $3$ is colourable.
Proof.
If $2\nmid|G|$ and $3\nmid|G|$, then the power maps $x\mapsto x^{2}$ and $x\mapsto x^{3}$ are bijections of $G$. Hence $\sigma(x)=x^{2}$ is a colouring bijection.
Thus, the question of colourability is particularly interesting for $2$-groups and $3$-groups. In this paper, we focus on $3$-groups.
We will make use of the above-mentioned result of A. B. Evans on the existence of strong complete mappings for noncyclic abelian $3$-groups.
Theorem 1.3.
([8]) Every noncyclic abelian $3$-group admits a strong complete mapping and hence is colourable.
The problem of lifting colouring bijections from quotient groups in the nonabelian case is more delicate. The lifting is straightforward in the case of quotients by central normal subgroups. The following lemma captures the mechanism that makes it possible to control the fibres over cosets.
Lemma 1.4 (Layer criterion).
Let $G$ be a group, $H\trianglelefteq G$, and let
$G=\bigsqcup_{t\in T}Ht$
be a decomposition into left cosets of $H$.
Assume that for each $i\in\{1,2,3\}$ and each $t\in T$ there exist a bijection
$F_{i}^{(t)}\colon H\to H$
and an element $t_{i}\in T$ such that for all $h\in H$
$\Delta^{(i)}_{\sigma}(ht)=F_{i}^{(t)}(h)\,t_{i}.$
If the maps $t\mapsto t_{i}$ are bijections of $T$, then each map $\Delta^{(i)}_{\sigma}\colon G\to G$ is bijective.
Proof.
For fixed $t\in T$, the map $\Delta^{(i)}_{\sigma}$ sends the coset $Ht$ onto $Ht_{i}$ via
$ht\mapsto F_{i}^{(t)}(h)t_{i}.$
Since $F_{i}^{(t)}$ is a bijection of $H$, this restriction is a bijection $Ht\to Ht_{i}$. Because $t\mapsto t_{i}$ is a bijection of $T$, the cosets $Ht_{i}$ run through all cosets of $H$ in $G$. Hence $\Delta^{(i)}_{\sigma}$ is a bijection of $G$. ∎
Lemma 1.5.
Let $G$ be a group, $H\trianglelefteq G$, and let $T\subseteq G$ be a transversal of $H$ in $G$. Let $\pi\colon G\to G/H$ be the natural projection and let $\tau\colon G/H\to T$ be the associated section.
Suppose that $\Phi\colon G/H\to G/H$ is a colouring bijection. Define
$\varphi\colon T\to T,\qquad\varphi(t):=\tau(\Phi(\pi(t))).$
For $t\in T$, writing $q=\pi(t)$, define
$t_{1}(t):=\tau(\Phi(q)q),\qquad t_{2}(t):=\tau(q^{-1}\Phi(q)),\qquad t_{3}(t):=\tau(q^{-1}\Phi(q)q).$
Then:
  1. (1)
    $\varphi\colon T\to T$ is a bijection;
  2. (2)
    for every $t\in T$ there exist uniquely determined elements $\xi(t),\zeta(t),\omega(t)\in H$ such that
    $\varphi(t)t=\xi(t)t_{1},\qquad t^{-1}\varphi(t)=\zeta(t)t_{2},\qquad t^{-1}\varphi(t)t=\omega(t)t_{3};$
  3. (3)
    the maps $\mapsto t_{1}, tt$$\mapsto t_{2}$, and $t\mapsto t_{3}$ are bijections of $T$.
Proof.
Since $\Phi$ is a bijection on $Q:=G/H$ and $\tau$ selects exactly one representative from each coset, the map $\varphi$ is a bijection of $T$.
Because $T$ is a transversal of $H$ in $G$, every element of $G$ admits a unique decomposition $ht$ with $h\in H$ and $t\in T$. Hence the three elements $\varphi(t)t$, $t^{-1}\varphi(t)$, and $t^{-1}\varphi(t)t$ determine uniquely the elements $\xi(t),\zeta(t),\omega(t)\in H$ and $t_{1},t_{2},t_{3}\in T$ appearing above.
Finally, since $\Phi$ is a colouring bijection on $Q$, the maps
$q\mapsto\Phi(q)q,\qquad q\mapsto q^{-1}\Phi(q),\qquad q\mapsto q^{-1}\Phi(q)q$
are bijections of $Q$. Identifying $T$ with $G/H$ via the bijection $\pi|_{T}:T\to G/H$, we may write
$t_{i}=\tau\bigl(\Delta_{\Phi}^{(i)}(\pi(t))\bigr).$
Since each $\Delta_{\Phi}^{(i)}$ is a bijection of $G/H$, it follows that the maps $t\mapsto t_{i}$ are bijections of $T$. ∎
Lemma 1.6.
Let $G$ be a group and let $H\leq Z(G)$ be a central subgroup. Suppose that $\Phi\colon G/H\to G/H$ is a colouring bijection and $\psi\colon H\to H$ is a colouring bijection of $H$. Then $G$ is colourable.
Proof.
Apply Lemma 1.5 to the colouring bijection $\Phi\colon G/H\to G/H$. Then for every $t\in T$ there exist uniquely determined $\xi(t),\zeta(t),\omega(t)\in H$ and $t_{1},t_{2},t_{3}\in T$ such that
$\varphi(t)t=\xi(t)t_{1},\qquad t^{-1}\varphi(t)=\zeta(t)t_{2},\qquad t^{-1}\varphi(t)t=\omega(t)t_{3},$
and the maps $t\mapsto t_{i}$ are bijections of $T$.
Define $\sigma\colon G\to G$ by
$\sigma(ht)=\psi(h)\varphi(t),\qquad h\in H,\ t\in T.$
Clearly $\sigma$ is a bijection.
Let $g=ht$ with $h\in H$ and $t\in T$. Using the centrality of $H$, we obtain
$\Delta^{(1)}_{\sigma}(ht)=(\psi(h)h)\,\varphi(t)t=(\psi(h)h\,\xi(t))\cdot t_{1},$
$\Delta^{(2)}_{\sigma}(ht)=(h^{-1}\psi(h))\,t^{-1}\varphi(t)=(h^{-1}\psi(h)\zeta(t))\cdot t_{2},$
$\Delta^{(3)}_{\sigma}(ht)=(h^{-1}\psi(h)h)\,t^{-1}\varphi(t)t=\psi(h)\omega(t)\cdot t_{3}.$
Thus each map $\Delta^{(i)}_{\sigma}$ has the form
$\Delta^{(i)}_{\sigma}(ht)=F^{(t)}_{i}(h)\,t_{i},$
where $F^{(t)}_{i}:H\to H$ are given explicitly by
$F^{(t)}_{1}(h)=\psi(h)h\,\xi(t),\qquad F^{(t)}_{2}(h)=h^{-1}\psi(h)\,\zeta(t),\qquad F^{(t)}_{3}(h)=\psi(h)\,\omega(t).$
Each $F^{(t)}_{i}$ is a bijection, being a composition of a bijection $\Delta_{\psi}^{(i)}$ with a right translation in $H$.
Therefore the maps $\Delta^{(i)}_{\sigma}$ are bijective by Lemma 1.4. Hence $\sigma$ is a colouring bijection of $G$. ∎
2. Colourability of groups of small order
Lemma 2.1.
All noncyclic groups of order at most $27$ are colourable.
Proof.
By Theorem 1.3, it suffices to consider the two nonabelian groups of order $27$: the Heisenberg group $H_{3}$ and the group
$L_{3}=\langle a,b\mid a^{9}=b^{3}=1,\ bab^{-1}=a^{4}\rangle.$
Recall that
$H_{3}=\langle x,y,z\mid x^{3}=y^{3}=z^{3}=1,\ xz=zx,\ yz=zy,\ xy=zyx\rangle.$
Every element of $H_{3}$ can be written uniquely in the form $x^{i}y^{j}z^{k}$ with $i,j,k\in\{0,1,2\}$.
We use coordinate notation for both groups. The permutations were obtained using the GAP system [13]. An exhaustive search was performed over permutations fixing the identity, and we verified whether the associated mapss $\Delta^{(1)}$, $\Delta^{(2)}$ and $\Delta^{(3)}$ were bijections.
The Heisenberg group. We identify elements of $H_{3}$ with triples $(i,j,k)\in\mathbb{Z}_{3}^{3}$, corresponding to $x^{i}y^{j}z^{k}$, with multiplication given by
$(i,j,k)\cdot(r,s,t)=(i+r,\;j+s,\;k+t+is),$
where all coordinates are taken modulo $3$.
The group $L_{3}$. We identify elements of $L_{3}$ with pairs $(i,j)\in\mathbb{Z}_{9}\times\mathbb{Z}_{3}$, corresponding to $a^{i}b^{j}$, with multiplication given by
$(i,j)\cdot(r,s)=(i+4^{j}r,\;j+s),$
where the first coordinate is taken modulo $9$ and the second modulo $3$.
Table 1. Colouring bijections for the groups $H_{3}$ and $L_{3}$ in additive notation.
Heisenberg group $H_{3}$ Group $L_{3}$
$x$ $\sigma(x)$ $\Delta^{(1)}(x)$ $\Delta^{(2)}(x)$ $\Delta^{(3)}(x)$ $x$ $\sigma(x)$ $\Delta^{(1)}(x)$ $\Delta^{(2)}(x)$ $\Delta^{(3)}(x)$
$(0,0,0)$ $(0,0,0)$ $(0,0,0)$ $(0,0,0)$ $(0,0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$
$(0,0,1)$ $(2,0,1)$ $(2,0,2)$ $(2,0,0)$ $(2,0,1)$ $(0,1)$ $(2,0)$ $(2,1)$ $(5,2)$ $(5,0)$
$(0,0,2)$ $(1,2,1)$ $(1,2,0)$ $(1,2,2)$ $(1,2,1)$ $(0,2)$ $(4,1)$ $(4,0)$ $(7,2)$ $(7,1)$
$(0,1,0)$ $(1,2,2)$ $(1,0,0)$ $(1,1,2)$ $(1,2,0)$ $(1,0)$ $(3,1)$ $(7,1)$ $(2,1)$ $(6,1)$
$(0,1,1)$ $(1,0,1)$ $(1,1,0)$ $(1,2,0)$ $(1,0,2)$ $(1,1)$ $(1,2)$ $(8,0)$ $(0,1)$ $(4,2)$
$(0,1,2)$ $(0,0,1)$ $(0,1,0)$ $(0,2,2)$ $(0,0,1)$ $(1,2)$ $(7,1)$ $(2,0)$ $(6,2)$ $(4,1)$
$(0,2,0)$ $(0,0,2)$ $(0,2,2)$ $(0,1,2)$ $(0,0,2)$ $(2,0)$ $(4,2)$ $(0,2)$ $(2,2)$ $(7,2)$
$(0,2,1)$ $(2,0,2)$ $(2,2,1)$ $(2,1,1)$ $(2,0,0)$ $(2,1)$ $(6,0)$ $(8,1)$ $(1,2)$ $(6,0)$
$(0,2,2)$ $(2,2,1)$ $(2,1,1)$ $(2,0,2)$ $(2,2,2)$ $(2,2)$ $(4,0)$ $(6,2)$ $(8,1)$ $(7,0)$
$(1,0,0)$ $(1,1,2)$ $(2,1,2)$ $(0,1,1)$ $(1,1,1)$ $(3,0)$ $(1,1)$ $(4,1)$ $(7,1)$ $(1,1)$
$(1,0,1)$ $(1,0,0)$ $(2,0,1)$ $(0,0,2)$ $(1,0,0)$ $(3,1)$ $(0,1)$ $(3,2)$ $(6,0)$ $(0,1)$
$(1,0,2)$ $(0,1,0)$ $(1,1,2)$ $(2,1,0)$ $(0,1,2)$ $(3,2)$ $(7,2)$ $(1,1)$ $(7,0)$ $(1,2)$
$(1,1,0)$ $(0,1,1)$ $(1,2,1)$ $(2,0,1)$ $(0,1,0)$ $(4,0)$ $(5,1)$ $(3,1)$ $(1,1)$ $(8,1)$
$(1,1,1)$ $(2,1,0)$ $(0,2,0)$ $(1,0,2)$ $(2,1,1)$ $(4,1)$ $(2,2)$ $(3,0)$ $(4,1)$ $(2,2)$
$(1,1,2)$ $(2,2,0)$ $(0,0,1)$ $(1,1,0)$ $(2,2,0)$ $(4,2)$ $(3,0)$ $(7,2)$ $(5,1)$ $(3,0)$
$(1,2,0)$ $(0,1,2)$ $(1,0,2)$ $(2,2,0)$ $(0,1,1)$ $(5,0)$ $(1,0)$ $(6,0)$ $(5,0)$ $(1,0)$
$(1,2,1)$ $(1,0,2)$ $(2,2,2)$ $(0,1,0)$ $(1,0,1)$ $(5,1)$ $(8,2)$ $(7,0)$ $(3,1)$ $(5,2)$
$(1,2,2)$ $(2,2,2)$ $(0,1,2)$ $(1,0,0)$ $(2,2,1)$ $(5,2)$ $(8,1)$ $(1,0)$ $(3,2)$ $(2,1)$
$(2,0,0)$ $(1,1,1)$ $(0,1,1)$ $(2,1,2)$ $(1,1,2)$ $(6,0)$ $(8,0)$ $(5,0)$ $(2,0)$ $(8,0)$
$(2,0,1)$ $(1,2,0)$ $(0,2,1)$ $(2,2,1)$ $(1,2,2)$ $(6,1)$ $(2,1)$ $(8,2)$ $(8,0)$ $(5,1)$
$(2,0,2)$ $(0,2,1)$ $(2,2,0)$ $(1,2,1)$ $(0,2,0)$ $(6,2)$ $(0,2)$ $(6,1)$ $(3,0)$ $(0,2)$
$(2,1,0)$ $(0,2,0)$ $(2,0,0)$ $(1,1,1)$ $(0,2,2)$ $(7,0)$ $(6,2)$ $(1,2)$ $(8,2)$ $(3,2)$
$(2,1,1)$ $(2,1,2)$ $(1,2,2)$ $(0,0,1)$ $(2,1,2)$ $(7,1)$ $(7,0)$ $(5,1)$ $(0,2)$ $(4,0)$
$(2,1,2)$ $(2,0,0)$ $(1,1,1)$ $(0,2,0)$ $(2,0,2)$ $(7,2)$ $(5,2)$ $(0,1)$ $(1,0)$ $(8,2)$
$(2,2,0)$ $(1,1,0)$ $(0,0,2)$ $(2,2,2)$ $(1,1,0)$ $(8,0)$ $(3,2)$ $(5,2)$ $(4,2)$ $(6,2)$
$(2,2,1)$ $(0,2,2)$ $(2,1,0)$ $(1,0,1)$ $(0,2,1)$ $(8,1)$ $(6,1)$ $(2,2)$ $(4,0)$ $(3,1)$
$(2,2,2)$ $(2,1,1)$ $(1,0,1)$ $(0,2,1)$ $(2,1,0)$ $(8,2)$ $(5,0)$ $(4,2)$ $(6,1)$ $(2,0)$
Remark 2.2.
Observe that the automorphism group $\operatorname{Aut}(G)$ acts naturally on the set of colouring bijections by conjugation. Thus, if $\varphi\in\operatorname{Aut}(G)$ and $\sigma$ is a colouring bijection, then
$\tau=\varphi^{-1}\circ\sigma\circ\varphi$
is also a colouring bijection. Indeed,
$\Delta_{\tau}^{(i)}=\varphi^{-1}\circ\Delta_{\sigma}^{(i)}\circ\varphi,\qquad i=1,2,3.$
For $H_{3}$ we have
$\operatorname{Aut}(H_{3})\cong(C_{2}\times C_{3})\rtimes\operatorname{GL}(2,\mathbb{F}_{3}),\qquad|\operatorname{Aut}(H_{3})|=432.$
A GAP computation shows that the stabiliser of $\sigma$ in $\operatorname{Aut}(H_{3})$ is trivial. Thus the orbit of $\sigma$ under $\operatorname{Aut}(H_{3})$ has size $432$. In particular, there are at least $432$ distinct colouring bijections of $H_{3}$.
Similarly,
$\operatorname{Aut}(L_{3})\cong(C_{3}\times C_{3})\rtimes C_{6}.$
Again, the stabiliser of $\sigma$ is trivial, and hence the orbit $\operatorname{Orb}(\sigma)$ contains $54$ distinct colouring bijections.
3. Lifting colouring bijections
We now prove two lifting results for colouring bijections from the quotient group $G/H$ to $G$ in the cases where $H\cong C_{3}\times C_{3}$ or $H\cong C_{9}\times C_{3}$. In the first case, we treat $H$ as a vector space over $\mathbb{F}_{3}$ and use certain automorphisms from the group $\operatorname{GL}(2,\mathbb{F}_{3})$. In the second case, we use a family of bijections of the group $C_{9}\times C_{3}$ determined computationally.
We begin with a simple lemma that will be useful in both situations.
Lemma 3.1.
Let $G$ be a $3$-group and suppose
$H=\langle c\rangle\times\langle b\rangle\trianglelefteq G,\qquad c\in Z(G),\qquad|b|=3^{k}$
for some $k\geq 1$. Then for every $t\in G$ there exists an integer
$m(t)\in\{0,1,\dots,3^{k-1}-1\}$
such that
$tbt^{-1}\in b^{\,1+3m(t)}\langle c\rangle.$
Equivalently, there exist integers $m(t)$ and $l(t)$ such that
$tbt^{-1}=b^{\,1+3m(t)}c^{\,l(t)}.$
If moreover $|c|=3$, then $l(t)\in\{0,1,2\}$ and, writing $z=b^{3}$, we obtain
$tbt^{-1}=b\,z^{m(t)}c^{l(t)}.$
Proof.
Fix $t\in G$. Since $H\trianglelefteq G$, we have $tbt^{-1}\in H$. Because $c\in Z(G)$, conjugation by $t$ fixes $c$ and therefore induces an automorphism of
$H/\langle c\rangle\cong\langle b\rangle\cong C_{3^{k}}.$
Thus we obtain a homomorphism
$G\to\operatorname{Aut}(C_{3^{k}})\cong(\mathbb{Z}/3^{k}\mathbb{Z})^{\times}.$
Since $G$ is a $3$-group, its image is a $3$-subgroup of $(\mathbb{Z}/3^{k}\mathbb{Z})^{\times}$. The unique Sylow $3$-subgroup of $(\mathbb{Z}/3^{k}\mathbb{Z})^{\times}$ is
{ $\{\,1+3m\pmod{3^{k}}:m=0,1,\dots,3^{k-1}-1\,\}.$
Hence for some $m(t)$ we have
$tbt^{-1}\equiv b^{\,1+3m(t)}\pmod{\langle c\rangle}$,
so
$tbt^{-1}=b^{\,1+3m(t)}c^{\,l(t)}$
for some integer $l(t)$. If $|c|=3$, then $l(t)$ may be chosen from $\{0,1,2\}$. ∎
3.1. Case $H=C_{3}\times C_{3}$
Consider the matrices in $\operatorname{GL}(2,\mathbb{F}_{3})$
$M_{1}=\begin{pmatrix}0&1\\ 1&1\end{pmatrix},\qquad M_{2}=\begin{pmatrix}0&2\\ 1&2\end{pmatrix}$.
For $\lambda\in\mathbb{F}_{3}$, define
$C(\lambda)=\begin{pmatrix}1&\lambda\\ 0&1\end{pmatrix}$.
Lemma 3.2.
For every $\lambda\in\mathbb{F}_{3}$ there exists $j\in\{1,2\}$ such that
$\det(M_{j}+I)\neq 0,\qquad\det(M_{j}-C(\lambda))\neq 0,\qquad\det(M_{j}+I-C(\lambda))\neq 0.$
Proof.
The determinants of the corresponding matrices are
$\det(M_{1}+I)=1,\qquad\det(M_{1}-C(\lambda))=\lambda-1,\qquad\det(M_{1}+I-C(\lambda))=\lambda-1,$
and
$\det(M_{2}+I)=1,\qquad\det(M_{2}-C(\lambda))=\lambda,\qquad\det(M_{2}+I-C(\lambda))=\lambda-2.$
Hence for $\lambda=0$ we have $\lambda-1\neq 0$, so all three matrices $M_{1}+I$, $M_{1}-C(\lambda)$, $M_{1}+I-C(\lambda)$ are invertible and we may take $M_{1}$. For $\lambda=1$ we have $\lambda\neq 0$ and $\lambda-2\neq 0$, so the three matrices corresponding to $M_{2}$ are invertible and we take $M_{2}$. For $\lambda=2$ we again have $\lambda-1\neq 0$, so we may take $M_{1}$.
In each case we obtain a matrix $M_{j}$ such that all three matrices
$M_{j}+I,\qquad M_{j}-C(\lambda),\qquad M_{j}+I-C(\lambda)$
are invertible. This completes the proof. ∎
Theorem 3.3.
Let $G$ be a $3$-group with a normal subgroup $H\cong C_{3}\times C_{3}$. If $G/H$ is colourable, then $G$ is also colourable.
Proof.
By Lemma 1.6, it suffices to consider the case where $H$ is not contained in the centre of $G$.
Suppose that $H\not\subseteq Z(G)$. Since $G$ is a $3$-group and $H\trianglelefteq G$ is nontrivial, we have $H\cap Z(G)\neq\{1\}$. By assumption, it follows that $|H\cap Z(G)|=3$.
Choose an element $z\in H\cap Z(G)$ of order $3$. Then $H/\langle z\rangle$ has order $3$, so we may choose $b\in H\setminus\langle z\rangle$. Since $H\cong C_{3}\times C_{3}$, every nontrivial element has order $3$, and therefore
$H=\langle z\rangle\times\langle b\rangle,\qquad z\in Z(G),\ b\notin Z(G).$
By Lemma 3.1, for any $g\in G$
$\rho_{g}(b)=gbg^{-1}=z^{k(g)}b.$
Every element of $H$ can be written uniquely as $z^{x}b^{y}$ with $x,y\in\mathbb{F}_{3}$, giving an identification
$H\cong\mathbb{F}_{3}^{2}.$
With respect to the basis corresponding to the generators $\{z,b\}$, automorphisms of $H$ are represented by matrices in $\operatorname{GL}(2,\mathbb{F}_{3})$. We will denote by $M(f)$ the matrix of the mapping $f$ in this basis.
For $t\in G$ the conjugation map $\rho_{t}$ acts on $H$ as an automorphism. Since $z\in Z(G)$,
$\rho_{t}(z)=z,\qquad\rho_{t}(b)=z^{k(t)}b,$
so
$M(\rho_{t})=C(k(t)).$
By Lemma 1.5, for each $t\in T$ there exist uniquely determined elements $\xi(t),\zeta(t),\omega(t)\in H$ and $t_{1},t_{2},t_{3}\in T$ such that
$\varphi(t)t=\xi(t)t_{1},\qquad t^{-1}\varphi(t)=\zeta(t)t_{2},\qquad t^{-1}\varphi(t)t=\omega(t)t_{3},$
and the maps $t\mapsto t_{i}$ are bijections of $T$.
For fixed $t$, define
$\alpha=k(\varphi(t)^{-1}).$
Then
$M(\rho_{\varphi(t)^{-1}})=C(\alpha),\qquad M(\rho_{\varphi(t)})=C(-\alpha).$
Choose $M_{t}\in\{M_{1},M_{2}\}$ according to Lemma 3.2. Define a linear automorphism $\psi_{t}\in\operatorname{Aut}(H)$ via the matrix
$M(\psi_{t})=M(\rho_{\varphi(t)})M_{t}.$
Finally, define
$\sigma(ht)=\psi_{t}(h)\varphi(t).$
We prove that $\sigma$ is a colouring bijection.
For $g=ht$ we have
$\Delta^{(1)}_{\sigma}(ht)$ $=\sigma(ht)\cdot(ht)=\bigl(\psi_{t}(h)\,\varphi(t)\bigr)\cdot(ht)=\psi_{t}(h)\cdot\bigl(\varphi(t)\,h\,\varphi(t)^{-1}\bigr)\cdot\bigl(\varphi(t)\,t\bigr)$
$=\psi_{t}(h)\cdot\rho_{\varphi(t)}(h)\cdot\xi(t)\,t_{1}.$
$\Delta^{(2)}_{\sigma}(ht)$ $=(ht)^{-1}\sigma(ht)=t^{-1}h^{-1}\cdot\psi_{t}(h)\,\varphi(t)=t^{-1}\cdot\bigl(h^{-1}\psi_{t}(h)\bigr)\cdot\varphi(t)$
$=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\cdot(t^{-1}\varphi(t))=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\cdot\zeta(t)\,t_{2}.$
$\Delta^{(3)}_{\sigma}(ht)$ $=\Delta^{(2)}_{\sigma}(ht)\cdot(ht)=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\cdot(t^{-1}\varphi(t))\cdot(ht)$
$=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\cdot\rho_{t^{-1}\varphi(t)}(h)\cdot(t^{-1}\varphi(t)t)$
$=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\cdot\rho_{t^{-1}\varphi(t)}(h)\cdot\omega(t)t_{3}.$
Define affine maps $F_{i}^{(t)}\colon H\to H$ by the $H$-components above. Their linear parts are given by the matrices
$M(F_{1}^{(t)})=M(\rho_{\varphi(t)})(M_{t}+I),$
$M(F_{2}^{(t)})=M(\rho_{t^{-1}\varphi(t)})(M_{t}-C(\alpha)),$
$M(F_{3}^{(t)})=M(\rho_{t^{-1}\varphi(t)})(M_{t}-C(\alpha)+I).$
By Lemma 3.2,
$M_{t}+I,\quad M_{t}-C(\alpha),\quad M_{t}-C(\alpha)+I$
are invertible, and matrices of the form $C(\cdot)$ are invertible. Hence all $F_{i}^{(t)}$ are bijections $H\to H$.
By Lemma 1.4, since the maps $t\mapsto t_{i}$ are bijections of the transversal $T$, each map $\Delta^{(i)}_{\sigma}\colon G\to G$ is bijective. As $\sigma$ itself is bijective, it is a colouring bijection of $G$. ∎
3.2. Case $H=C_{3}\times C_{9}$
We prove the existence of a lift of a colouring bijection of the $3$-group $G/H$ to a colouring bijection of $G$ in the case $H\cong C_{3}\times C_{9}$. The situation is asymmetric, since an element of order $9$ may or may not lie in the centre of $G$. This affects the form of the conjugations $\rho_{x}$ acting on $H$, and therefore the proof requires a separate treatment of these two cases.
Lemma 3.4.
Let
$H=\langle c\rangle\times\langle b\rangle\cong C_{9}\times C_{3},$
where $|b|=3$ and $|c|=9$. Set $z:=c^{3}$.
There exist bijections $\alpha_{0},\alpha_{1},\alpha_{2}\colon H\to H$ such that:
  • the maps $\Delta^{(1)}_{\alpha_{0}}$ and $\Delta^{(2)}_{\alpha_{0}}$ are bijections;
  • for $\lambda=1,2$, both $\Delta^{(2)}_{\alpha_{\lambda}}$ and the map
    $T_{\lambda}\colon H\to H,\qquad T_{\lambda}(h)=z^{\lambda j}\,h\,\alpha_{\lambda}(h),\text{ where }h=c^{u}b^{j},$
    are bijections.
Proof.
Identify
$H=\langle c\rangle\times\langle b\rangle\cong\mathbb{Z}_{9}\times\mathbb{Z}_{3}$
via $c^{u}b^{j}\leftrightarrow(u,j)$. Then $z=c^{3}=(3,0)$ and the group operation is componentwise addition.
Let $\alpha_{0},\alpha_{1},\alpha_{2}\colon H\to H$ be the maps given in Table LABEL:Tab:3-add. In additive notation, for $h\in H$,
$\Delta^{(1)}_{\alpha_{0}}(h)=h+\alpha_{0}(h),\qquad\Delta^{(2)}_{\alpha_{\lambda}}(h)=-h+\alpha_{\lambda}(h),$
and
$T_{\lambda}(h)=\lambda j\,z+h+\alpha_{\lambda}(h),\qquad\text{for }h=(u,j),\ \lambda=1,2.$
Direct inspection of Table LABEL:Tab:3-add, shows that each listed column is a permutation of $H$. Hence all these maps are bijections. ∎
Lemma 3.5.
Let
$H=\langle b\rangle\times\langle c\rangle\cong C_{9}\times C_{3},$
where $|b|=9$, $|c|=3$, and set $z=b^{3}$.
For $\lambda,\ell\in\{0,1,2\}$ define
$q_{\lambda,\ell}(b^{j}c^{m})=z^{\lambda j}c^{\ell j}.$
Then for every pair $(\lambda,\ell)\in\{0,1,2\}^{2}$ there exists a bijection
$f_{\lambda,\ell}\colon H\to H$
such that the three maps
$A_{\lambda,\ell}(h):=q_{\lambda,\ell}(h)\,h\,f_{\lambda,\ell}(h),\qquad B_{\lambda,\ell}(h):=h^{-1}f_{\lambda,\ell}(h),\qquad C_{\lambda,\ell}(h):=q_{\lambda,\ell}(h)\,f_{\lambda,\ell}(h)$
are bijections of $H$.
Proof.
Write $H=\mathbb{Z}_{9}\times\mathbb{Z}_{3}$ additively, identifying $b^{j}c^{m}$ with $(j,m)$. Then
$q_{\lambda,\ell}(j,m)=(3\lambda j,\ell j).$
For each $(\lambda,\ell)\in\{0,1,2\}^{2}$, a bijection $f_{\lambda,\ell}\colon H\to H$ is given explicitly in Appendix B. Using these maps, define
$A_{\lambda,\ell}(h)=q_{\lambda,\ell}(h)+h+f_{\lambda,\ell}(h),\quad B_{\lambda,\ell}(h)=-h+f_{\lambda,\ell}(h),\quad C_{\lambda,\ell}(h)=q_{\lambda,\ell}(h)+f_{\lambda,\ell}(h).$
By direct inspection of the tables in Appendix B, each of the maps $f_{\lambda,\ell}$, $A_{\lambda,\ell}$, $B_{\lambda,\ell}$, and $C_{\lambda,\ell}$ is a permutation of $H$. Hence all these maps are bijections, which proves the lemma. ∎
Theorem 3.6.
Let $G$ be a finite $3$-group and let $H\cong C_{3}\times C_{9}$ be an abelian normal subgroup such that the quotient group $G/H$ is colourable. Then $G$ is also colourable.
Proof.
Since $C_{3}\times C_{9}$ is colourable (see Theorem 1.3), by Lemma 1.6 it suffices to consider the case where $H$ is not contained in the centre of $G$.
Because $H$ contains central elements, we may assume that
$H=\langle c,b\rangle,\qquad c\in Z(G).$
Then either $|c|=9$ and $|b|=3$, or $|c|=3$ and $|b|=9$.
We apply Lemma 1.5. Let $\Phi\colon G/H\to G/H$ be a colouring bijection and define
$\varphi(t)=\tau(\Phi(\pi(t))).$
Thus, for each $t\in T$ there exist uniquely determined elements $\xi(t),\zeta(t),\omega(t)\in H$ and $t_{1},t_{2},t_{3}\in T$ such that
$\varphi(t)t=\xi(t)t_{1},\qquad t^{-1}\varphi(t)=\zeta(t)t_{2},\qquad t^{-1}\varphi(t)t=\omega(t)t_{3},$
and the maps $t\mapsto t_{i}$ are bijections of $T$.
Case 1: $|c|=9$ and $|b|=3$.
For $x\in G$, denote by
$\rho_{x}\colon H\to H,\qquad\rho_{x}(h)=xhx^{-1}$
the conjugation automorphism of $H$ induced by $x$.
By Lemma 3.1, for every $x\in G$ we have
$\rho_{x}(b)=z^{k(x)}b,\qquad z=c^{3},\qquad k(x)\in\{0,1,2\}.$
Moreover, for any $h\in H$ and $t\in T$,
$\rho_{ht}(b)=(ht)b(ht)^{-1}=h(tbt^{-1})h^{-1}=hz^{k(t)}bh^{-1}=z^{k(t)}b=\rho_{t}(b).$
Define
$u(t):=k\bigl(\varphi(t)\bigr),\qquad u(t)\in\{0,1,2\}.$
Define a map $\sigma\colon G\to G$ by
$\sigma(ht)=\alpha_{u(t)}(h)\,\varphi(t).$
We now determine the explicit form of conjugation on $H$. For $x\in G$ and $h=b^{j}c^{m}\in H$ we have
$\rho_{x}(b^{j}c^{m})=z^{k(x)j}b^{j}c^{m}.$
Indeed, since $H$ is abelian and $z\in Z(G)$,
$\rho_{x}(b^{j}c^{m})=(xbx^{-1})^{j}(xcx^{-1})^{m}=(z^{k(x)}b)^{j}c^{m}=z^{k(x)j}b^{j}c^{m}.$
Since $u(t)=k(\varphi(t))$, for $h=b^{j}c^{m}$ we obtain
($\ast$) $\rho_{\varphi(t)}(h)=z^{u(t)j}h.$
Let $g=ht$ with $h=b^{j}c^{m}\in H$ and $t\in T$. Using $\varphi(t)h=\rho_{\varphi(t)}(h)\varphi(t)$ we compute
$\Delta^{(1)}_{\sigma}(ht)=\bigl(\alpha_{u(t)}(h)\,z^{u(t)j}h\bigr)(\varphi(t)t)=T_{u(t)}(h)(\varphi(t)t)=T_{u(t)}(h)\xi(t)\cdot t_{1}$
where for $u(t)=0$ we set
$T_{0}=\Delta^{(1)}_{\alpha_{0}}$.
Similarly, we obtain
$\Delta^{(2)}_{\sigma}(ht)=\rho_{t^{-1}}\!\bigl(\Delta^{(2)}_{\alpha_{u(t)}}(h)\bigr)\,(t^{-1}\varphi(t))=\rho_{t^{-1}}\!\bigl(\Delta^{(2)}_{\alpha_{u(t)}}(h)\bigr)\,\zeta(t)\cdot t_{2},$
and
$\Delta^{(3)}_{\sigma}(ht)=\rho_{t^{-1}}\!\bigl(T_{u(t)}(h)\bigr)\,(t^{-1}\varphi(t)t)=\rho_{t^{-1}}\!\bigl(T_{u(t)}(h)\bigr)\,\omega(t)\cdot t_{3}.$
Thus
$\Delta^{(1)}_{\sigma}(ht)=F_{1}^{(t)}(h)t_{1},$
$\Delta^{(2)}_{\sigma}(ht)=F_{2}^{(t)}(h)t_{2},$
$\Delta^{(3)}_{\sigma}(ht)=F_{3}^{(t)}(h)t_{3}.$
For each fixed $t\in T$, the map $F_{1}^{(t)}$ is a translation composed with $T_{u(t)}$, the map $F_{2}^{(t)}$ is a translation composed with $\rho_{t^{-1}}\circ\Delta^{(2)}_{\alpha_{u(t)}}$, and the map $F_{3}^{(t)}$ is a translation composed with $\rho_{t^{-1}}\circ T_{u(t)}$. Hence Lemma 3.4 implies that all maps $F_{i}^{(t)}\colon H\to H$ are bijections. By Lemma 1.5, the maps $t\mapsto t_{i}$ are bijections of $T$, and by Lemma 1.4, each map $\Delta^{(i)}_{\sigma}$ is bijective, and therefore $\sigma$ is a colouring bijection.
Case 2: $|c|=3$ and $|b|=9$, where $c\in Z(G)$.
By Lemma 3.1, for every $x\in G$ there exist $k(x),l(x)\in\{0,1,2\}$ such that
$\rho_{x}(b)=b^{1+3k(x)}c^{l(x)}.$
Since $c\in Z(G)$ and $z=b^{3}$, it follows that for every $h=b^{j}c^{m}\in H$,
$\rho_{x}(h)=\rho_{x}(b^{j}c^{m})=z^{k(x)j}c^{l(x)j}b^{j}c^{m}.$
Set
$\lambda(t)=k(\varphi(t)),\qquad\ell(t)=l(\varphi(t)).$
We will use Lemma 3.5. Recall that
$q_{\lambda,\ell}(b^{j}c^{m})=z^{\lambda j}c^{\ell j}.$
Define
$\sigma(ht)=f_{\lambda(t),\ell(t)}(h)\varphi(t),\qquad h\in H,\ t\in T.$
Since each $f_{\lambda(t),\ell(t)}$ is a bijection of $H$ and $\varphi\colon T\to T$ is a bijection, the map $\sigma\colon G\to G$ is bijective.
Let $g=ht$ with $h=b^{j}c^{m}\in H$ and $t\in T$. Since
$\rho_{\varphi(t)}(h)=q_{\lambda(t),\ell(t)}(h)\,h,$
we obtain
$\Delta^{(1)}_{\sigma}(ht)$ $=\sigma(ht)\,ht=f_{\lambda(t),\ell(t)}(h)\,\varphi(t)\,h\,t$
$=f_{\lambda(t),\ell(t)}(h)\,\rho_{\varphi(t)}(h)\,(\varphi(t)t)$
$=f_{\lambda(t),\ell(t)}(h)\,q_{\lambda(t),\ell(t)}(h)\,h\,(\varphi(t)t).$
Since $H$ is abelian,
$f_{\lambda(t),\ell(t)}(h)\,q_{\lambda(t),\ell(t)}(h)\,h=q_{\lambda(t),\ell(t)}(h)\,h\,f_{\lambda(t),\ell(t)}(h)=A_{\lambda(t),\ell(t)}(h).$
Hence
$\Delta^{(1)}_{\sigma}(ht)=A_{\lambda(t),\ell(t)}(h)\,\xi(t)\cdot t_{1}.$
Similarly,
$\Delta^{(2)}_{\sigma}(ht)$ $=(ht)^{-1}\sigma(ht)=t^{-1}h^{-1}f_{\lambda(t),\ell(t)}(h)\varphi(t)$
$=\rho_{t^{-1}}\!\bigl(h^{-1}f_{\lambda(t),\ell(t)}(h)\bigr)\,(t^{-1}\varphi(t))$
$=\rho_{t^{-1}}\!\bigl(B_{\lambda(t),\ell(t)}(h)\bigr)\,\zeta(t)\cdot t_{2}.$
Finally,
$\Delta^{(3)}_{\sigma}(ht)$ $=(ht)^{-1}\sigma(ht)(ht)=t^{-1}h^{-1}f_{\lambda(t),\ell(t)}(h)\varphi(t)ht$
$=t^{-1}h^{-1}f_{\lambda(t),\ell(t)}(h)\rho_{\varphi(t)}(h)\,\varphi(t)t$
$=t^{-1}h^{-1}f_{\lambda(t),\ell(t)}(h)q_{\lambda(t),\ell(t)}(h)h\,\varphi(t)t.$
Again using that $H$ is abelian, we get
$h^{-1}f_{\lambda(t),\ell(t)}(h)q_{\lambda(t),\ell(t)}(h)h=q_{\lambda(t),\ell(t)}(h)f_{\lambda(t),\ell(t)}(h)=C_{\lambda(t),\ell(t)}(h).$
Therefore
$\Delta^{(3)}_{\sigma}(ht)=\rho_{t^{-1}}\!\bigl(C_{\lambda(t),\ell(t)}(h)\bigr)\,\omega(t)\cdot t_{3}.$
Thus the maps
$F_{1}^{(t)}(h):=A_{\lambda(t),\ell(t)}(h)\,\xi(t),\qquad F_{2}^{(t)}(h):=\rho_{t^{-1}}\!\bigl(B_{\lambda(t),\ell(t)}(h)\bigr)\,\zeta(t),$
$F_{3}^{(t)}(h):=\rho_{t^{-1}}\!\bigl(C_{\lambda(t),\ell(t)}(h)\bigr)\,\omega(t)$
are bijections of $H$. This follows from Lemma 3.5, because $A_{\lambda(t),\ell(t)}$, $B_{\lambda(t),\ell(t)}$, and $C_{\lambda(t),\ell(t)}$ are bijections, while $\rho_{t^{-1}}$ and right multiplication by $\xi(t),\zeta(t),\omega(t)$ are bijections of $H$. Since $\Phi$ is a colouring bijection on $G/H$, the maps $t\mapsto t_{i}$ are bijections of $T$. Hence, by Lemma 1.4, each of the maps
$\Delta^{(1)}_{\sigma},\qquad\Delta^{(2)}_{\sigma},\qquad\Delta^{(3)}_{\sigma}$
is bijective. Hence $\sigma$ is a colouring bijection of $G$. ∎
4. Main theorem
The aim of this section is to prove the main theorem. A crucial role is played by the following structural result describing $3$-groups with a cyclic maximal subgroup.
Theorem 4.1.
(cf. [11, Theorem 5.3.4]) Let $G$ be a $3$-group of order $3^{r}$ having a cyclic maximal subgroup. Then either
$G\cong C_{3^{r-1}}\times C_{3}\quad\text{or}\quad G\cong L_{r}.$
The following auxiliary lemma will be used later. Since we could not find a suitable reference, we include a short proof for completeness. We write $\Phi(G)$ for the Frattini subgroup of $G$ and $\Omega_{1}(G)=\langle x\in G\mid x^{3}=1\rangle$.
Lemma 4.2.
Let $G$ be a $3$-group of order greater than $27$. If every maximal subgroup of $G$ is noncyclic, then $G$ contains a normal subgroup $H$ isomorphic to $C_{3}\times C_{3}$ and such that $G/H$ is not cyclic.
Proof.
First observe that if a $3$-group $G$ contains a normal elementary abelian subgroup $H$ of rank at least $2$, then $G$ contains a normal subgroup $N$ of rank $2$, and hence $N\cong C_{3}\times C_{3}$. Indeed, the conjugation action of $G$ on $H$ induces a homomorphism
$G\to\operatorname{GL}(H).$
Its image is a $3$-group. Over the field $\mathbb{F}_{3}$, every element of $3$-power order is unipotent, so the image acts unipotently on $H$. Hence $H$ admits a $G$-invariant flag
$0<H_{1}<H_{2}<\cdots<H.$
Since $\dim_{\mathbb{F}_{3}}H\geq 2$, the subgroup $H_{2}$ has order $9$ and is normal in $G$. Thus $H_{2}\cong C_{3}\times C_{3}$.
Let $V=\Omega_{1}(Z(\Phi(G)))$. We show that $V$ is noncyclic. Suppose to the contrary that it is, and choose $G$ of minimal order subject to the hypotheses. Since $V\leq Z(\Phi(G))$, the subgroup $V$ is characteristic in $\Phi(G)$ and hence normal in $G$. As $V$ is cyclic of order $3$, $\operatorname{Aut}(V)\cong C_{2}$, and therefore the conjugation action of the $3$-group $G$ on $V$ is trivial. Hence $V\leq Z(G)$.
Put $\overline{G}=G/V$. Because $V\leq\Phi(G)$, we have
$\Phi(\overline{G})=\Phi(G)/V.$
We claim that every maximal subgroup of $\overline{G}$ is noncyclic. Indeed, let $\overline{M}=M/V$ be a maximal subgroup of $\overline{G}$. If $\overline{M}$ were cyclic, then $M/V$ would be cyclic. Since $V\leq\Phi(G)\leq\Phi(M)$, the group $M/\Phi(M)$ would be cyclic. By Burnside’s basis theorem, this implies $d(M)=1$, so $M$ is cyclic, contrary to the hypothesis.
If $|\overline{G}|>27$, then by the minimality of $G$ the group $\overline{G}$ contains a normal subgroup $\overline{N}\cong C_{3}\times C_{3}$. Let $N$ be its full preimage in $G$. Then $N\lhd G$, $|N|=27$, and $N/V\cong C_{3}\times C_{3}$. Since $V\leq Z(N)$ and $|V|=3$, the classification of groups of order $27$ shows that $N$ contains a characteristic subgroup $E\cong C_{3}\times C_{3}$. Hence $E\lhd G$, a contradiction. Thus $|\overline{G}|=27$.
A group of order $27$ whose maximal subgroups are all noncyclic is either $C_{3}^{3}$ or the Heisenberg group $H_{3}$. In either case, $\overline{G}$ contains a normal subgroup $\overline{N}\cong C_{3}\times C_{3}$. Let $N$ be its full preimage in $G$. As above, $N$ contains a characteristic subgroup $E\cong C_{3}\times C_{3}$, which is therefore normal in $G$. This contradiction shows that $V$ is noncyclic.
Thus $V$ is an elementary abelian normal subgroup of rank at least $2$. By the observation made at the beginning of the proof, $G$ contains a normal subgroup isomorphic to $C_{3}\times C_{3}$. Since $\dim_{\mathbb{F}_{3}}(G/\Phi(G))\geq 2$ and $H\leq\Phi(G)$, the factor group $G/H$ is noncyclic. ∎
We now prove the main theorem.
Theorem 4.3.
Every noncyclic $3$-group $G$ not isomorphic to $L_{r}$ for $r\geq 4$ is colourable.
Proof.
Suppose that the theorem is false, and let $G$ be a counterexample of minimal order. Write $|G|=3^{r}$. By Lemma 2.1, every noncyclic group of order at most $27$ is colourable, hence $r\geq 4$.
Assume first that $G$ has a cyclic maximal subgroup. Then by Theorem 4.1
$G\cong C_{3^{r-1}}\times C_{3}\quad\text{or}\quad G\cong L_{r}.$
In the first case, $G$ is colourable by Theorem 1.3, while the second contradicts the hypothesis of the theorem. Hence $G$ has no cyclic maximal subgroup.
By Lemma 4.2, the group $G$ contains a normal subgroup $N\cong C_{3}\times C_{3}$ such that $G/N$ is not cyclic.
By the minimality of $G$, either $G/N$ is colourable or
$G/N\cong L_{r-2}.$
If $G/N$ is colourable, then Theorem 3.3 implies that $G$ is colourable, a contradiction. Hence
$G/N\cong L_{r-2}.$
Since $L_{3}$ is colourable by Lemma 2.1, we must have $r\geq 6$.
Choose generators $z,b\in N$ such that
$N=\langle z,b\rangle,\qquad z\in Z(G),\qquad|z|=|b|=3.$
Let $\bar{x},\bar{y}$ be generators of $G/N\cong L_{r-2}$ satisfying
$\bar{x}^{3^{r-3}}=\bar{y}^{3}=1,\qquad\bar{y}\bar{x}\bar{y}^{-1}=\bar{x}^{1+3^{r-4}}.$
Choose lifts $x,y\in G$ with $xN=\bar{x}$ and $yN=\bar{y}$. Then
$G=\langle x,y,z,b\rangle.$
By Lemma 3.1, for each $t\in G$ there exists $k(t)\in\{0,1,2\}$ such that
$tbt^{-1}=z^{k(t)}b.$
Since $z$ is central and $|z|=3$, we obtain
$b^{-1}tb=z^{k(t)}t,\qquad b^{-1}t^{3}b=t^{3},$
so $b,t^{3}]=1$ for all $t\in G$. Set
$c=x^{3^{r-4}}.$
Since $x$ has order $3^{r-3}$ modulo $N$, the element $c=x^{3^{r-4}}$ has order $3$ and lies in the centre of $L_{r-2}$. In particular, $c$ commutes with $x$. Then $[b,c]=1$. Moreover, from
$\bar{y}\bar{x}\bar{y}^{-1}=\bar{x}^{1+3^{r-4}}=\bar{c}\,\bar{x}$
we obtain
$yxy^{-1}=z^{i}b^{j}cx$
for some $i,j\in\{0,1,2\}$. Because $z$ is central, $[b,c]=1$, and $c$ commutes with $x$, we have
$yx^{3}y^{-1}$ $=(yxy^{-1})^{3}=(z^{i}b^{j}cx)^{3}=z^{3i}\,(b^{j}cx)^{3}$
$=(b^{j}c)(xb^{j}cx^{-1})(x^{2}b^{j}cx^{-2})x^{3}.$
Now $xbx^{-1}=z^{k(x)}b$, hence
$x^{\nu}b^{j}x^{-\nu}=z^{\nu jk(x)}b^{j},\quad(\nu=0,1,2).$
Using also $z^{3}=1$ and $(b^{j})^{3}=1$, we obtain
$yx^{3}y^{-1}=c^{3}x^{3}.$
As $c^{3}\in N$ and $N$ has exponent $3$, we have $c^{9}=1$. Therefore
$ycy^{-1}=yx^{3^{r-4}}y^{-1}=(yx^{3}y^{-1})^{3^{r-5}}=(c^{3}x^{3})^{3^{r-5}}=c^{3^{r-4}}x^{3^{r-4}}=c.$
Clearly $c$ commutes with $x$, $z$, and $b$, so $c\in Z(G)$. Furthermore, $c\notin N$, since otherwise $G/N$ would be abelian, contradicting $G/N\cong L_{r-2}$. Now let
$S=\langle c,z\rangle\leq Z(G).$
Case 1. $|c|=3$. Then $S\cong C_{3}\times C_{3}$ and
$yx^{3}y^{-1}=x^{3},$
so $x^{3}\in Z(G)$. Put
$d=x^{3^{r-5}}.$
Since $r\geq 6$, we have $d\in Z(G)$ and $d^{3}=c$. Moreover, for every $t\in G$ we have
$tbt^{-1}=z^{k(t)}b\equiv b\pmod{S}$,
so $bS\in Z(G/S)$. Also $dS\in Z(G/S)$ because $d\in Z(G)$.
Now $dS$ and $bS$ both have order $3$ in $G/S$. We claim that
$\langle dS\rangle\neq\langle bS\rangle.$
Indeed, otherwise $dS=b^{\varepsilon}S$ for some $\varepsilon\in\{1,2\}$, and therefore
$d=b^{\varepsilon}c^{u}z^{v}$
for suitable integers $u,v$. Passing modulo $N=\langle z,b\rangle$, we obtain
$dN=c^{u}N.$
But $c=x^{3^{r-4}}\in\langle x\rangle$, so $cN$ has order $3$ in $G/N\cong L_{r-2}$, whereas
$d=x^{3^{r-5}}$
implies that $dN$ has order $9$. This contradiction proves the claim.
Hence $Z(G/S)$ contains at least two distinct subgroups of order $3$, so $Z(G/S)$ is not cyclic. Since the group $L_{r-2}$ has cyclic centre of order $3$, it follows that
$G/S\not\cong L_{r-2}.$
The quotient $G/S$ is not cyclic and is not isomorphic to any group $L_{k}$, hence by the minimality of $G$ it is colourable. Applying Lemma 1.6, we conclude that $G$ is colourable, a contradiction.
Case 2. $|c|=9$. Then $c^{3}\in N$, so we may write
$c^{3}=z^{i}b^{j}$
with $i,j\in\{0,1,2\}$.
Subcase 2a: $j\neq 0$. Then $b\in\langle c,z\rangle$, and hence
$S=\langle c,z\rangle\cong C_{9}\times C_{3}\qquad\text{and}\qquad N=\langle z,b\rangle\leq S.$
Moreover,
$yxy^{-1}=z^{i}b^{j}cx\equiv x\pmod{S}$,
so $G/S$ is abelian.
Since $N\leq S$, we have
$G/S\cong(G/N)/(S/N).$
Now $G/N\cong L_{r-2}$, and $S/N$ is generated by $cN=x^{3^{r-4}}N$, which is the central subgroup of order $3$ in $L_{r-2}$. Therefore
$G/S\cong L_{r-2}/Z(L_{r-2})\cong C_{3^{\,r-4}}\times C_{3}.$
In particular, $G/S$ is abelian and noncyclic. Hence $G/S$ is colourable by Theorem 1.3. Since $S\leq Z(G)$, Lemma 1.6 implies that $G$ is colourable, a contradiction.
Subcase 2b: $j=0$. Then
$c^{3}=z^{i},$
so
$H=\langle c,b\rangle\cong C_{9}\times C_{3}.$
Since $c\in Z(G)$ and $[b,c]=1$, the subgroup $H$ is abelian and normal in $G$.
We now show that $G/H$ is abelian and noncyclic. First,
$yxy^{-1}=z^{i}cx=c^{3}cx=c^{4}x\equiv x\pmod{H}$,
so $G/H$ is abelian.
Next, since $N=\langle z,b\rangle\leq H$, we have
$G/H\cong(G/N)/(H/N).$
Now $H/N$ is generated by $cN$, and $cN=x^{3^{r-4}}N$ is precisely the central subgroup of order $3$ in $G/N\cong L_{r-2}$. Hence
$G/H\cong L_{r-2}/Z(L_{r-2})\cong C_{3^{\,r-4}}\times C_{3}.$
Thus $G/H$ is abelian and noncyclic. By Theorem 1.3, both the quotient $G/H$ and the subgroup $H\cong C_{9}\times C_{3}$ are colourable, and by Theorem 3.6 $G$ is colourable, again a contradiction.
All cases lead to a contradiction, completing the proof. ∎
Concluding remarks
  1. (1)
    Every finite nilpotent group is a direct product of its Sylow subgroups. Hence, if $G$ has odd order and its Sylow $3$-subgroup is noncyclic and not isomorphic to any $L_{r}$ ($r\geq 4$), then $G$ is colourable. In particular:
    Corollary 4.4.
    Let $G$ be a nilpotent group of odd order whose Sylow $3$-subgroup is neither cyclic of order greater than $1$ nor isomorphic to any group $L_{r}$ ($r\geq 4$). Then
    $\chi(\mathscr{G}_{3}(G))=|G|.$
  2. (2)
    The colourability of the groups $L_{r}$ for $r\geq 4$ remains open. On the other hand, if $G$ is any nontrivial $3$-group, then $L_{r}\times G$ is colourable, since its centre is noncyclic and the product is therefore not isomorphic to any $L_{s}$.
  3. (3)
    By Theorem 4.3, every $3$-group without a cyclic subgroup of index $3$ is colourable. A natural next problem is the case of $2$-groups. It is plausible that the absence of cyclic subgroups of index $4$ is sufficient; the recent results of [2] on strong complete mappings support this expectation. Such a theorem would settle the colourability problem for finite nilpotent groups.
  4. (4)
    Another natural direction is to study the colourability of solvable groups, particularly those of order $2^{m}3^{n}$.
  5. (5)
    For nonabelian groups, the three defining bijectivity conditions are genuinely independent: there exist bijections $\sigma$ for which two of the maps $\Delta^{(i)}_{\sigma}$ are bijective, while the third is not.
    In particular, for the group $L_{3}$ there exists a permutation $\sigma$ such that $\Delta^{(1)}_{\sigma}$ and $\Delta^{(2)}_{\sigma}$ are bijections, whereas $\Delta^{(3)}_{\sigma}$ is not. It is given by
    $\sigma=$ $(1\;ab\;a^{2}b^{2}\;a^{3}b\;a^{2}b\;a^{7}\;a^{7}b^{2}\;a\;a^{8}b^{2}\;a^{5}b\;a^{8}\;a^{6}\;a^{6}b\;a^{5}\;ab^{2}\;a^{2}\;a^{8}b\;a^{4}b^{2}\;a^{3})$
    $(b\;a^{4}b\;a^{3}b^{2}\;a^{5}b^{2}\;a^{6}b^{2}\;b^{2}\;a^{7}b),$
    where $a^{4}$ is a fixed point. This shows that, in the nonabelian setting, the three bijectivity conditions in the definition are not redundant.
Appendix A Bijections for $\mathbb{Z}_{9}\times\mathbb{Z}_{3}$
The permutations were obtained computationally using the GAP system [13].
Table 2. Bijections for Lemma 3.4
$h$ $\alpha_{0}(h)$ $\Delta^{(1)}_{\alpha_{0}}(h)$ $\Delta^{(2)}_{\alpha_{0}}(h)$ $\alpha_{1}(h)$ $\Delta^{(2)}_{\alpha_{1}}(h)$ $\alpha_{2}(h)$ $\Delta^{(2)}_{\alpha_{2}}(h)$ $T_{1}(h)$ $T_{2}(h)$
$(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$
$(1,0)$ $(2,0)$ $(3,0)$ $(1,0)$ $(2,0)$ $(1,0)$ $(2,0)$ $(1,0)$ $(3,0)$ $(3,0)$
$(2,0)$ $(7,1)$ $(0,1)$ $(5,1)$ $(7,2)$ $(5,2)$ $(2,1)$ $(0,1)$ $(0,2)$ $(4,1)$
$(3,0)$ $(2,1)$ $(5,1)$ $(8,1)$ $(2,2)$ $(8,2)$ $(0,2)$ $(6,2)$ $(5,2)$ $(3,2)$
$(4,0)$ $(6,1)$ $(1,1)$ $(2,1)$ $(3,1)$ $(8,1)$ $(8,1)$ $(4,1)$ $(7,1)$ $(3,1)$
$(5,0)$ $(1,0)$ $(6,0)$ $(5,0)$ $(1,0)$ $(5,0)$ $(1,0)$ $(5,0)$ $(6,0)$ $(6,0)$
$(6,0)$ $(8,2)$ $(5,2)$ $(2,2)$ $(4,2)$ $(7,2)$ $(8,2)$ $(2,2)$ $(1,2)$ $(5,2)$
$(7,0)$ $(2,2)$ $(0,2)$ $(4,2)$ $(1,1)$ $(3,1)$ $(3,1)$ $(5,1)$ $(8,1)$ $(1,1)$
$(8,0)$ $(7,2)$ $(6,2)$ $(8,2)$ $(4,1)$ $(5,1)$ $(7,2)$ $(8,2)$ $(3,1)$ $(6,2)$
$(0,1)$ $(3,0)$ $(3,1)$ $(3,2)$ $(3,0)$ $(3,2)$ $(3,0)$ $(3,2)$ $(6,1)$ $(0,1)$
$(1,1)$ $(3,1)$ $(4,2)$ $(2,0)$ $(1,2)$ $(0,1)$ $(8,0)$ $(7,2)$ $(5,0)$ $(6,1)$
$(2,1)$ $(8,1)$ $(1,2)$ $(6,0)$ $(8,1)$ $(6,0)$ $(1,1)$ $(8,0)$ $(4,2)$ $(0,2)$
$(3,1)$ $(8,0)$ $(2,1)$ $(5,2)$ $(7,0)$ $(4,2)$ $(2,2)$ $(8,1)$ $(4,1)$ $(2,0)$
$(4,1)$ $(4,0)$ $(8,1)$ $(0,2)$ $(4,0)$ $(0,2)$ $(4,0)$ $(0,2)$ $(2,1)$ $(5,1)$
$(5,1)$ $(5,2)$ $(1,0)$ $(0,1)$ $(0,1)$ $(4,0)$ $(0,1)$ $(4,0)$ $(8,2)$ $(2,2)$
$(6,1)$ $(1,2)$ $(7,0)$ $(4,1)$ $(8,2)$ $(2,1)$ $(4,2)$ $(7,1)$ $(8,0)$ $(7,0)$
$(7,1)$ $(5,1)$ $(3,2)$ $(7,0)$ $(5,1)$ $(7,0)$ $(4,1)$ $(6,0)$ $(6,2)$ $(8,2)$
$(8,1)$ $(6,2)$ $(5,0)$ $(7,1)$ $(0,2)$ $(1,1)$ $(5,2)$ $(6,1)$ $(2,0)$ $(1,0)$
$(0,2)$ $(4,2)$ $(4,1)$ $(4,0)$ $(3,2)$ $(3,0)$ $(5,1)$ $(5,2)$ $(0,1)$ $(8,0)$
$(1,2)$ $(7,0)$ $(8,2)$ $(6,1)$ $(5,0)$ $(4,1)$ $(3,2)$ $(2,0)$ $(3,2)$ $(7,1)$
$(2,2)$ $(5,0)$ $(7,2)$ $(3,1)$ $(8,0)$ $(6,1)$ $(5,0)$ $(3,1)$ $(7,2)$ $(1,2)$
$(3,2)$ $(1,1)$ $(4,0)$ $(7,2)$ $(5,2)$ $(2,0)$ $(7,1)$ $(4,2)$ $(5,1)$ $(4,0)$
$(4,2)$ $(3,2)$ $(7,1)$ $(8,0)$ $(6,1)$ $(2,2)$ $(6,0)$ $(2,1)$ $(7,0)$ $(4,2)$
$(5,2)$ $(6,0)$ $(2,2)$ $(1,1)$ $(2,1)$ $(6,2)$ $(6,1)$ $(1,2)$ $(4,0)$ $(5,0)$
$(6,2)$ $(0,2)$ $(6,1)$ $(3,0)$ $(7,1)$ $(1,2)$ $(7,0)$ $(1,1)$ $(1,0)$ $(7,2)$
$(7,2)$ $(4,1)$ $(2,0)$ $(6,2)$ $(6,2)$ $(8,0)$ $(1,2)$ $(3,0)$ $(1,1)$ $(2,1)$
$(8,2)$ $(0,1)$ $(8,0)$ $(1,2)$ $(6,0)$ $(7,1)$ $(6,2)$ $(7,0)$ $(2,2)$ $(8,1)$
Appendix B Tables for Lemma 3.5
In all tables, the coordinate pair $(j,m)$ denotes the element $b^{j}c^{m}\in H\cong C_{9}\times C_{3}$.
Table 3. The maps $f_{\lambda,\ell}$ for $H\cong C_{9}\times C_{3}$
$\lambda=0$ $\lambda=1$ $\lambda=2$
$h$ $\ell=0$ $\ell=1$ $\ell=2$ $\ell=0$ $\ell=1$ $\ell=2$ $\ell=0$ $\ell=1$ $\ell=2$
$(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$
$(0,1)$ $(1,0)$ $(1,0)$ $(1,2)$ $(1,0)$ $(1,1)$ $(1,0)$ $(7,2)$ $(1,1)$ $(6,2)$
$(0,2)$ $(2,0)$ $(5,2)$ $(5,0)$ $(5,2)$ $(2,0)$ $(2,2)$ $(4,0)$ $(2,2)$ $(4,0)$
$(1,0)$ $(5,2)$ $(1,1)$ $(1,1)$ $(7,2)$ $(8,1)$ $(4,0)$ $(0,1)$ $(8,0)$ $(0,1)$
$(1,1)$ $(8,0)$ $(7,1)$ $(7,1)$ $(4,2)$ $(4,0)$ $(5,1)$ $(4,2)$ $(7,1)$ $(2,1)$
$(1,2)$ $(6,0)$ $(6,1)$ $(3,0)$ $(3,1)$ $(3,2)$ $(7,2)$ $(5,2)$ $(4,2)$ $(6,0)$
$(2,0)$ $(6,1)$ $(5,1)$ $(5,1)$ $(4,1)$ $(6,1)$ $(5,2)$ $(7,0)$ $(7,0)$ $(0,2)$
$(2,1)$ $(0,2)$ $(2,0)$ $(2,0)$ $(1,1)$ $(7,2)$ $(8,0)$ $(3,2)$ $(1,2)$ $(5,0)$
$(2,2)$ $(3,2)$ $(3,2)$ $(6,0)$ $(6,1)$ $(1,0)$ $(7,0)$ $(0,2)$ $(6,2)$ $(7,1)$
$(3,0)$ $(4,1)$ $(7,0)$ $(7,2)$ $(7,0)$ $(7,0)$ $(2,1)$ $(5,1)$ $(5,2)$ $(4,1)$
$(3,1)$ $(1,1)$ $(0,2)$ $(0,2)$ $(0,1)$ $(4,2)$ $(4,2)$ $(4,1)$ $(4,0)$ $(7,0)$
$(3,2)$ $(2,1)$ $(4,0)$ $(4,2)$ $(4,0)$ $(0,1)$ $(6,0)$ $(8,0)$ $(3,1)$ $(5,1)$
$(4,0)$ $(4,2)$ $(8,1)$ $(8,0)$ $(2,1)$ $(7,1)$ $(1,1)$ $(1,1)$ $(0,2)$ $(4,2)$
$(4,1)$ $(8,1)$ $(2,2)$ $(2,1)$ $(8,2)$ $(1,2)$ $(0,1)$ $(5,0)$ $(8,2)$ $(1,0)$
$(4,2)$ $(7,2)$ $(6,2)$ $(3,2)$ $(3,0)$ $(3,1)$ $(3,2)$ $(3,1)$ $(2,0)$ $(3,2)$
$(5,0)$ $(7,0)$ $(8,2)$ $(8,2)$ $(5,1)$ $(5,2)$ $(1,2)$ $(1,2)$ $(8,1)$ $(3,0)$
$(5,1)$ $(2,2)$ $(4,2)$ $(4,0)$ $(2,2)$ $(6,0)$ $(3,0)$ $(7,1)$ $(4,1)$ $(7,2)$
$(5,2)$ $(5,0)$ $(3,1)$ $(6,1)$ $(6,2)$ $(2,2)$ $(4,1)$ $(2,2)$ $(5,0)$ $(8,2)$
$(6,0)$ $(8,2)$ $(2,1)$ $(2,2)$ $(0,2)$ $(5,0)$ $(6,2)$ $(8,2)$ $(2,1)$ $(5,2)$
$(6,1)$ $(3,1)$ $(0,1)$ $(0,1)$ $(8,1)$ $(8,0)$ $(7,1)$ $(3,0)$ $(7,2)$ $(3,1)$
$(6,2)$ $(0,1)$ $(8,0)$ $(8,1)$ $(2,0)$ $(0,2)$ $(8,1)$ $(6,0)$ $(3,0)$ $(2,2)$
$(7,0)$ $(3,0)$ $(5,0)$ $(5,2)$ $(3,2)$ $(3,0)$ $(5,0)$ $(1,0)$ $(0,1)$ $(2,0)$
$(7,1)$ $(1,2)$ $(6,0)$ $(3,1)$ $(5,0)$ $(5,1)$ $(0,2)$ $(6,1)$ $(1,0)$ $(1,2)$
$(7,2)$ $(6,2)$ $(4,1)$ $(4,1)$ $(1,2)$ $(2,1)$ $(2,0)$ $(2,1)$ $(6,1)$ $(8,1)$
$(8,0)$ $(7,1)$ $(1,2)$ $(1,0)$ $(6,0)$ $(6,2)$ $(6,1)$ $(6,2)$ $(3,2)$ $(6,1)$
$(8,1)$ $(4,0)$ $(3,0)$ $(6,2)$ $(8,0)$ $(8,2)$ $(8,2)$ $(2,0)$ $(6,0)$ $(1,1)$
$(8,2)$ $(5,1)$ $(7,2)$ $(7,0)$ $(7,1)$ $(4,1)$ $(3,1)$ $(8,1)$ $(5,1)$ $(8,0)$
Table 4. The maps $A_{\lambda,\ell}$ for $H\cong C_{9}\times C_{3}$ in coordinates $(j,m)$ corresponding to $b^{j}c^{m}$.
$\lambda=0$ $\lambda=1$ $\lambda=2$
$h$ $\ell=0$ $\ell=1$ $\ell=2$ $\ell=0$ $\ell=1$ $\ell=2$ $\ell=0$ $\ell=1$ $\ell=2$
$(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$
$(0,1)$ $(1,1)$ $(1,1)$ $(1,0)$ $(1,1)$ $(1,2)$ $(1,1)$ $(7,0)$ $(1,2)$ $(6,0)$
$(0,2)$ $(2,2)$ $(5,1)$ $(5,2)$ $(5,1)$ $(2,2)$ $(2,1)$ $(4,2)$ $(2,1)$ $(4,2)$
$(1,0)$ $(6,2)$ $(2,2)$ $(2,0)$ $(2,2)$ $(3,2)$ $(8,2)$ $(7,1)$ $(6,1)$ $(7,0)$
$(1,1)$ $(0,1)$ $(8,0)$ $(8,1)$ $(8,0)$ $(8,2)$ $(0,1)$ $(2,0)$ $(5,0)$ $(0,1)$
$(1,2)$ $(7,2)$ $(7,1)$ $(4,1)$ $(7,0)$ $(7,2)$ $(2,0)$ $(3,1)$ $(2,2)$ $(4,1)$
$(2,0)$ $(8,1)$ $(7,0)$ $(7,2)$ $(3,1)$ $(5,0)$ $(4,0)$ $(3,0)$ $(3,2)$ $(5,0)$
$(2,1)$ $(2,0)$ $(4,0)$ $(4,2)$ $(0,2)$ $(6,2)$ $(7,2)$ $(8,0)$ $(6,2)$ $(1,2)$
$(2,2)$ $(5,1)$ $(5,0)$ $(8,0)$ $(5,0)$ $(0,1)$ $(6,0)$ $(5,1)$ $(2,0)$ $(3,1)$
$(3,0)$ $(7,1)$ $(1,0)$ $(1,2)$ $(1,0)$ $(1,0)$ $(5,1)$ $(8,1)$ $(8,2)$ $(7,1)$
$(3,1)$ $(4,2)$ $(3,0)$ $(3,0)$ $(3,2)$ $(7,0)$ $(7,0)$ $(7,2)$ $(7,1)$ $(1,1)$
$(3,2)$ $(5,0)$ $(7,2)$ $(7,1)$ $(7,2)$ $(3,0)$ $(0,2)$ $(2,2)$ $(6,0)$ $(8,0)$
$(4,0)$ $(8,2)$ $(3,2)$ $(3,2)$ $(0,1)$ $(5,2)$ $(8,0)$ $(2,1)$ $(1,0)$ $(5,1)$
$(4,1)$ $(3,2)$ $(6,1)$ $(6,1)$ $(6,0)$ $(8,1)$ $(7,1)$ $(6,1)$ $(0,1)$ $(2,0)$
$(4,2)$ $(2,1)$ $(1,2)$ $(7,0)$ $(1,2)$ $(1,1)$ $(1,0)$ $(4,0)$ $(3,0)$ $(4,0)$
$(5,0)$ $(3,0)$ $(4,1)$ $(4,0)$ $(7,1)$ $(7,1)$ $(3,0)$ $(0,2)$ $(7,0)$ $(2,1)$
$(5,1)$ $(7,0)$ $(0,2)$ $(0,2)$ $(4,0)$ $(8,0)$ $(5,2)$ $(6,2)$ $(3,1)$ $(6,1)$
$(5,2)$ $(1,2)$ $(8,2)$ $(2,1)$ $(8,1)$ $(4,0)$ $(6,1)$ $(1,1)$ $(4,1)$ $(7,2)$
$(6,0)$ $(5,2)$ $(8,1)$ $(8,2)$ $(6,2)$ $(2,0)$ $(3,2)$ $(5,2)$ $(8,1)$ $(2,2)$
$(6,1)$ $(0,2)$ $(6,2)$ $(6,2)$ $(5,2)$ $(5,1)$ $(4,2)$ $(0,1)$ $(4,0)$ $(0,2)$
$(6,2)$ $(6,0)$ $(5,2)$ $(5,0)$ $(8,2)$ $(6,1)$ $(5,0)$ $(3,2)$ $(0,2)$ $(8,1)$
$(7,0)$ $(1,0)$ $(3,1)$ $(3,1)$ $(4,2)$ $(4,1)$ $(6,2)$ $(5,0)$ $(4,2)$ $(6,2)$
$(7,1)$ $(8,0)$ $(4,2)$ $(1,1)$ $(6,1)$ $(6,0)$ $(1,2)$ $(1,2)$ $(5,2)$ $(5,2)$
$(7,2)$ $(4,1)$ $(2,1)$ $(2,2)$ $(2,1)$ $(3,1)$ $(3,1)$ $(6,0)$ $(1,1)$ $(3,2)$
$(8,0)$ $(6,1)$ $(0,1)$ $(0,1)$ $(2,0)$ $(2,1)$ $(2,2)$ $(8,2)$ $(5,1)$ $(8,2)$
$(8,1)$ $(3,1)$ $(2,0)$ $(5,1)$ $(4,1)$ $(4,2)$ $(4,1)$ $(4,1)$ $(8,0)$ $(3,0)$
$(8,2)$ $(4,0)$ $(6,0)$ $(6,0)$ $(3,0)$ $(0,2)$ $(8,1)$ $(1,0)$ $(7,2)$ $(1,0)$
Table 5. The maps $B_{\lambda,\ell}$ for $H\cong C_{9}\times C_{3}$ in coordinates $(j,m)$ corresponding to $b^{j}c^{m}$.
$\lambda=0$ $\lambda=1$ $\lambda=2$
$h$ $\ell=0$ $\ell=1$ $\ell=2$ $\ell=0$ $\ell=1$ $\ell=2$ $\ell=0$ $\ell=1$ $\ell=2$
$(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$
$(0,1)$ $(1,2)$ $(1,2)$ $(1,1)$ $(1,2)$ $(1,0)$ $(1,2)$ $(7,1)$ $(1,0)$ $(6,1)$
$(0,2)$ $(2,1)$ $(5,0)$ $(5,1)$ $(5,0)$ $(2,1)$ $(2,0)$ $(4,1)$ $(2,0)$ $(4,1)$
$(1,0)$ $(4,2)$ $(0,1)$ $(0,1)$ $(6,2)$ $(7,1)$ $(3,0)$ $(8,1)$ $(7,0)$ $(8,1)$
$(1,1)$ $(7,2)$ $(6,0)$ $(6,0)$ $(3,1)$ $(3,2)$ $(4,0)$ $(3,1)$ $(6,0)$ $(1,0)$
$(1,2)$ $(5,1)$ $(5,2)$ $(2,1)$ $(2,2)$ $(2,0)$ $(6,0)$ $(4,0)$ $(3,0)$ $(5,1)$
$(2,0)$ $(4,1)$ $(3,1)$ $(3,1)$ $(2,1)$ $(4,1)$ $(3,2)$ $(5,0)$ $(5,0)$ $(7,2)$
$(2,1)$ $(7,1)$ $(0,2)$ $(0,2)$ $(8,0)$ $(5,1)$ $(6,2)$ $(1,1)$ $(8,1)$ $(3,2)$
$(2,2)$ $(1,0)$ $(1,0)$ $(4,1)$ $(4,2)$ $(8,1)$ $(5,1)$ $(7,0)$ $(4,0)$ $(5,2)$
$(3,0)$ $(1,1)$ $(4,0)$ $(4,2)$ $(4,0)$ $(4,0)$ $(8,1)$ $(2,1)$ $(2,2)$ $(1,1)$
$(3,1)$ $(7,0)$ $(6,1)$ $(6,1)$ $(6,0)$ $(1,1)$ $(1,1)$ $(1,0)$ $(1,2)$ $(4,2)$
$(3,2)$ $(8,2)$ $(1,1)$ $(1,0)$ $(1,1)$ $(6,2)$ $(3,1)$ $(5,1)$ $(0,2)$ $(2,2)$
$(4,0)$ $(0,2)$ $(4,1)$ $(4,0)$ $(7,1)$ $(3,1)$ $(6,1)$ $(6,1)$ $(5,2)$ $(0,2)$
$(4,1)$ $(4,0)$ $(7,1)$ $(7,0)$ $(4,1)$ $(6,1)$ $(5,0)$ $(1,2)$ $(4,1)$ $(6,2)$
$(4,2)$ $(3,0)$ $(2,0)$ $(8,0)$ $(8,1)$ $(8,2)$ $(8,0)$ $(8,2)$ $(7,1)$ $(8,0)$
$(5,0)$ $(2,0)$ $(3,2)$ $(3,2)$ $(0,1)$ $(0,2)$ $(5,2)$ $(5,2)$ $(3,1)$ $(7,0)$
$(5,1)$ $(6,1)$ $(8,1)$ $(8,2)$ $(6,1)$ $(1,2)$ $(7,2)$ $(2,0)$ $(8,0)$ $(2,1)$
$(5,2)$ $(0,1)$ $(7,2)$ $(1,2)$ $(1,0)$ $(6,0)$ $(8,2)$ $(6,0)$ $(0,1)$ $(3,0)$
$(6,0)$ $(2,2)$ $(5,1)$ $(5,2)$ $(3,2)$ $(8,0)$ $(0,2)$ $(2,2)$ $(5,1)$ $(8,2)$
$(6,1)$ $(6,0)$ $(3,0)$ $(3,0)$ $(2,0)$ $(2,2)$ $(1,0)$ $(6,2)$ $(1,1)$ $(6,0)$
$(6,2)$ $(3,2)$ $(2,1)$ $(2,2)$ $(5,1)$ $(3,0)$ $(2,2)$ $(0,1)$ $(6,1)$ $(5,0)$
$(7,0)$ $(5,0)$ $(7,0)$ $(7,2)$ $(5,2)$ $(5,0)$ $(7,0)$ $(3,0)$ $(2,1)$ $(4,0)$
$(7,1)$ $(3,1)$ $(8,2)$ $(5,0)$ $(7,2)$ $(7,0)$ $(2,1)$ $(8,0)$ $(3,2)$ $(3,1)$
$(7,2)$ $(8,0)$ $(6,2)$ $(6,2)$ $(3,0)$ $(4,2)$ $(4,1)$ $(4,2)$ $(8,2)$ $(1,2)$
$(8,0)$ $(8,1)$ $(2,2)$ $(2,0)$ $(7,0)$ $(7,2)$ $(7,1)$ $(7,2)$ $(4,2)$ $(7,1)$
$(8,1)$ $(5,2)$ $(4,2)$ $(7,1)$ $(0,2)$ $(0,1)$ $(0,1)$ $(3,2)$ $(7,2)$ $(2,0)$
$(8,2)$ $(6,2)$ $(8,0)$ $(8,1)$ $(8,2)$ $(5,2)$ $(4,2)$ $(0,2)$ $(6,2)$ $(0,1)$
Table 6. The maps $C_{\lambda,\ell}$ for $H\cong C_{9}\times C_{3}$ in coordinates $(j,m)$ corresponding to $b^{j}c^{m}$.
$\lambda=0$ $\lambda=1$ $\lambda=2$
$h$ $\ell=0$ $\ell=1$ $\ell=2$ $\ell=0$ $\ell=1$ $\ell=2$ $\ell=0$ $\ell=1$ $\ell=2$
$(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$ $(0,0)$
$(0,1)$ $(1,0)$ $(1,0)$ $(1,2)$ $(1,0)$ $(1,1)$ $(1,0)$ $(7,2)$ $(1,1)$ $(6,2)$
$(0,2)$ $(2,0)$ $(5,2)$ $(5,0)$ $(5,2)$ $(2,0)$ $(2,2)$ $(4,0)$ $(2,2)$ $(4,0)$
$(1,0)$ $(5,2)$ $(1,2)$ $(1,0)$ $(1,2)$ $(2,2)$ $(7,2)$ $(6,1)$ $(5,1)$ $(6,0)$
$(1,1)$ $(8,0)$ $(7,2)$ $(7,0)$ $(7,2)$ $(7,1)$ $(8,0)$ $(1,2)$ $(4,2)$ $(8,0)$
$(1,2)$ $(6,0)$ $(6,2)$ $(3,2)$ $(6,1)$ $(6,0)$ $(1,1)$ $(2,2)$ $(1,0)$ $(3,2)$
$(2,0)$ $(6,1)$ $(5,0)$ $(5,2)$ $(1,1)$ $(3,0)$ $(2,0)$ $(1,0)$ $(1,2)$ $(3,0)$
$(2,1)$ $(0,2)$ $(2,2)$ $(2,1)$ $(7,1)$ $(4,1)$ $(5,1)$ $(6,2)$ $(4,1)$ $(8,1)$
$(2,2)$ $(3,2)$ $(3,1)$ $(6,1)$ $(3,1)$ $(7,2)$ $(4,1)$ $(3,2)$ $(0,1)$ $(1,2)$
$(3,0)$ $(4,1)$ $(7,0)$ $(7,2)$ $(7,0)$ $(7,0)$ $(2,1)$ $(5,1)$ $(5,2)$ $(4,1)$
$(3,1)$ $(1,1)$ $(0,2)$ $(0,2)$ $(0,1)$ $(4,2)$ $(4,2)$ $(4,1)$ $(4,0)$ $(7,0)$
$(3,2)$ $(2,1)$ $(4,0)$ $(4,2)$ $(4,0)$ $(0,1)$ $(6,0)$ $(8,0)$ $(3,1)$ $(5,1)$
$(4,0)$ $(4,2)$ $(8,2)$ $(8,2)$ $(5,1)$ $(1,2)$ $(4,0)$ $(7,1)$ $(6,0)$ $(1,1)$
$(4,1)$ $(8,1)$ $(2,0)$ $(2,0)$ $(2,2)$ $(4,0)$ $(3,0)$ $(2,0)$ $(5,0)$ $(7,2)$
$(4,2)$ $(7,2)$ $(6,0)$ $(3,1)$ $(6,0)$ $(6,2)$ $(6,1)$ $(0,1)$ $(8,1)$ $(0,1)$
$(5,0)$ $(7,0)$ $(8,1)$ $(8,0)$ $(2,1)$ $(2,1)$ $(7,0)$ $(4,2)$ $(2,0)$ $(6,1)$
$(5,1)$ $(2,2)$ $(4,1)$ $(4,1)$ $(8,2)$ $(3,2)$ $(0,1)$ $(1,1)$ $(7,0)$ $(1,0)$
$(5,2)$ $(5,0)$ $(3,0)$ $(6,2)$ $(3,2)$ $(8,1)$ $(1,2)$ $(5,2)$ $(8,2)$ $(2,0)$
$(6,0)$ $(8,2)$ $(2,1)$ $(2,2)$ $(0,2)$ $(5,0)$ $(6,2)$ $(8,2)$ $(2,1)$ $(5,2)$
$(6,1)$ $(3,1)$ $(0,1)$ $(0,1)$ $(8,1)$ $(8,0)$ $(7,1)$ $(3,0)$ $(7,2)$ $(3,1)$
$(6,2)$ $(0,1)$ $(8,0)$ $(8,1)$ $(2,0)$ $(0,2)$ $(8,1)$ $(6,0)$ $(3,0)$ $(2,2)$
$(7,0)$ $(3,0)$ $(5,1)$ $(5,1)$ $(6,2)$ $(6,1)$ $(8,2)$ $(7,0)$ $(6,2)$ $(8,2)$
$(7,1)$ $(1,2)$ $(6,1)$ $(3,0)$ $(8,0)$ $(8,2)$ $(3,1)$ $(3,1)$ $(7,1)$ $(7,1)$
$(7,2)$ $(6,2)$ $(4,2)$ $(4,0)$ $(4,2)$ $(5,2)$ $(5,2)$ $(8,1)$ $(3,2)$ $(5,0)$
$(8,0)$ $(7,1)$ $(1,1)$ $(1,1)$ $(3,0)$ $(3,1)$ $(3,2)$ $(0,2)$ $(6,1)$ $(0,2)$
$(8,1)$ $(4,0)$ $(3,2)$ $(6,0)$ $(5,0)$ $(5,1)$ $(5,0)$ $(5,0)$ $(0,2)$ $(4,2)$
$(8,2)$ $(5,1)$ $(7,1)$ $(7,1)$ $(4,1)$ $(1,0)$ $(0,2)$ $(2,1)$ $(8,0)$ $(2,1)$